# Multiple dispersive bounds. II) Sub-threshold branch-cuts

### v2

@@ VI.1 Unitary filtering of the input data
Starting from the original covariance matrix $C_{ij}$ for the input data, the action of the filter (74) is to select the unitary subset $\{\overline{f}_{i}\}$ with a covariance matrix $\overline{C}_{ij}$, i.e. only those data that can be reproduced exactly by (an infinite number of) unitary BGL expansions, given by Eq. (12) with the constraint (16). When the filter (74) is not fulfilled, then the input data cannot be reproduced exactly by any unitary BGL expansion. These data contain non-unitary effects and they are removed from the analysis (see for details Ref. [1]). Therefore, to the subset of input data satisfying the global filter (74) we apply the expansion (31), adopting the single total bound (34) for the coefficients of the expansion.
In the case of the analysis employing the double bound (32)-(33) we modify the selection of the unitary input data as follows. Since $\chi[\beta]$ is the integral over the unit circle given in Eq. (VI.1), we can split such an integral into the sum of two (positive) contributions, related to the pair-production arc and outside the pair-production arc, namely
@@
$\chi_{\rm pair}[\beta]$ $=$ $\sum_{i,j=1}^{N_{data}}\phi(z_{i})B(z_{i})d_{i}f_{i}\,\phi(z_{j})B(z_{j})d_{j}f_{j}\,\frac{(1-z_{i}^{2})(1-z_{j}^{2})}{1-z_{i}z_{j}}\left[1-B_{ij}(\alpha_{th})\right]~,$~ (78)
@@
In the kaon case of interest in this paper there are no poles below the lowest threshold $t_{th}=4m_{\pi}^{2}\simeq 0.08$ GeV2 and, therefore, $B(z)=1$. We apply either the global filter (74) or the double filter (81)-(82). The main effect of such filters is a modification of the original covariance matrix $C_{ij}$ into a new matrix $\overline{C}_{ij}$ by introducing appropriate correlations either between the two experimental datasets or among the LQCD data. We stress that such modifications are dictated by unitarity. It turns out that in the case of the experimental data from FNAL [6] and CERN [7] the impact of the filtering procedure described above on the predicted bands of values for the form factor is limited, while this is not the case when the LQCD data from the HotQCD Collaboration [8] are considered. This is likely due to the facts that correlations are not provided in Ref. [8] and the LQCD points exhibit a remarkable high precision (at the level of few permille). Consequently, the global filter (74) or the double filter (81)-(82) produce a new highly correlated covariance matrix $\overline{C}_{ij}$. The new predicted bands for the form factor $F_{K^{\pm}}^{(em)}(Q^{2})$ are shown in Fig. 9 and the impact of the filtering procedures is clearly visible by a direct comparison with Figs. 6 and 8.
@@
The anomalous increase of the width of the predicted bands at low $Q^{2}$, visible in Fig. 8, is completely eliminated. Moreover, at large $Q^{2}$ the width of the predicted bands is reduced significantly both at $q=0$ and $q=3.5$. Note that the predictions of the double bound expansion are nicely consistent with the LQCD points up to $Q^{2}\simeq 28$ GeV2 both at $q=0$ and $q=3.5$, while those corresponding to the total bound expansion are only marginally consistent. Finally, the most precise predictions at large $Q^{2}$ are those corresponding to the double bound expansion with the choice $q=0$.
VI.2 Charged kaon radius
@@
where the second error comes from the small dependence upon the choice of the kinematical function (estimated using our results obtained at $q=0$ and $q=3.5$). Thus, the value quoted by PDG underestimates the uncertainty by a factor of $\simeq 2$ due to the model-dependent assumption about the momentum dependence of the charged kaon form factor.
@@ VII Conclusions
We have discussed a specific application of the framework of multiple dispersive bounds, firstly introduced in the companion paper [1], to the study of sub-threshold branch-cuts. The main results are as follows.
  • In the standard BGL $z$-expansion we have proposed a double dispersive bound as the proper way to deal with the presence of sub-threshold branch-cuts in accordance with unitarity. The key idea is that two different unitarity constraints have to be considered at the same time, namely the first one in the the pair-production arc and the second one in the extra region between the lowest sub-threshold and the pair-production branch-points. We have highlighted that the bound related to the extra region, $\chi_{\rm extra}^{U}$, is not related to the usual susceptibilities, i.e. to the quantities coming from the computation of the derivatives of appropriate two-point Green functions in momentum space. It remains an open issue to understand the proper Euclidean correlation functions that would give direct access to the bound $\chi_{\rm extra}^{U}$ directly from first-principles, e.g. from lattice simulations.
@@
  • We have analyzed the experimental data or the lattice QCD (LQCD) results available for the electromagnetic form factor of the charged kaon in the spacelike-region. The comparison with other methodologies present in literature and with the BGL $z$-expansion based on the single, total dispersive bound has clearly shown that the $z$-expansion including the double dispersive bound provides the most precise extrapolation at large momentum transfer as well as the most stable results with respect to the choice of the outer function outside the pair-production region.
  • We have highlighted the impact of the application of the unitarity filter on the given set of input data (see Section VI.1 and Ref. [1]) and we have compared our model-independent findings for the charged kaon radius with the model-dependent estimates made in the last PDG review [9] and in Ref. [8] (see Section VI.2).

### v3

@@ VI.2 Unitary filtering of the input data
Starting from the original covariance matrix $C_{ij}$ for the input data and choosing the kinematical function $\phi(z)$ as in Eq. (12), the action of the filter (74) is to select the unitary subset $\{\overline{f}_{i}\}$ with a covariance matrix $\overline{C}_{ij}$, i.e. only those data that can be reproduced exactly by (an infinite number of) unitary BGL expansions, given by Eq. (12) with the constraint (16). When the filter (74) is not fulfilled, then the input data cannot be reproduced exactly by any unitary BGL expansion. These data contain non-unitary effects and they are removed from the analysis (see for details Ref. [1]). Therefore, to the subset of input data satisfying the global filter (74) we apply the expansion (31), adopting the single total bound (34) for the coefficients of the expansion.
In the case of the analysis employing the double bound (32)-(33) we modify the selection of the unitary input data as follows. Since $\chi[\beta]$ is the integral over the unit circle given in Eq. (VI.2), we can split such an integral into the sum of two (positive) contributions, related to the pair-production arc and outside the pair-production arc, namely
@@
$\chi_{\rm extra}[\beta]$ $=$ $\sum_{i,j=1}^{N_{data}}\phi(z_{i})B(z_{i})d_{i}f_{i}\,\phi(z_{j})B(z_{j})d_{j}f_{j}\,\frac{(1-z_{i}^{2})(1-z_{j}^{2})}{1-z_{i}z_{j}}\left[1-B_{ij}(\alpha_{th})\right]~,$~ (78)
@@
In the kaon case of interest in this paper there are no poles below the lowest threshold $t_{th}=4m_{\pi}^{2}\simeq 0.08$ GeV2 and, therefore, $B(z)=1$. We apply either the global filter (74) or the double filter (81)-(82). The main effect of such filters is a modification of the original covariance matrix $C_{ij}$ into a new matrix $\overline{C}_{ij}$ by introducing appropriate correlations either between the two experimental datasets or among the LQCD data. We stress that such modifications are dictated by unitarity. It turns out that in the case of the experimental data from FNAL [13] and CERN [14] the impact of the filtering procedure described above on the predicted bands of values for the form factor is limited, while this is not the case when the LQCD data from the HotQCD Collaboration [15] are considered. This is likely due to the fact that the LQCD results exhibit a remarkable high precision (at the level of few permille) and also that correlations are not provided at all in Ref. [15]. The global filter (74) as well as the double filter (81)-(82), that both depend on the choice of the kinematical function, may change the mean values and errors of the form factor and may produce a new highly correlated covariance matrix $\overline{C}_{ij}$ sensitive to the value of qq [^14] We have checked that the subset of filtered data is always well consistent with a normal probability distribution.. In order to give a quantitative estimate of the impact of the filtering procedure we have computed the values of the quantities $\Delta$ and $\epsilon$, defined in Eqs. (32)-(33) of the companion paper [1]. For both choices of $q$ and in the case of either the single or the double dispersive bound we always find $\Delta\simeq 0.6$ and $\epsilon\simeq 0.7$, which guarantee that anomalous modifications of mean values and errors of the form factor are excluded. The corresponding predicted bands for the form factor $F_{K^{\pm}}^{(em)}(Q^{2})$ are shown in Fig. 9 in the case of the LQCD data and the impact of the filtering procedures is clearly visible by a direct comparison with Figs. 6 and 8.
@@
In particular, though partially blurred by the size of the markers, the widths of the bands at low $Q^{2}$ in Fig. 8 ($q=3.5$) are definitely larger than the errors of the data points used in the fitting procedure, while the same does not occur in Fig. 6 ($q=0$). These findings are related to the $q$-dependent impact of non-unitary events in the (unfiltered) sample of input data. Indeed, as shown in Fig. 9, when the fitting procedure is applied only to the unitarity-filtered subset of input data, the widths of the bands at low $Q^{2}$ do not exceed the errors of the data points for both values of $q$. Moreover, at large $Q^{2}$ the width of the predicted bands is reduced significantly both at $q=0$ and $q=3.5$.
We point out that the predictions of either the total or the double bound expansion are consistent with the LQCD points above $Q^{2}\simeq 0.4$ GeV2 up to $Q^{2}\simeq 28$ GeV2 both at $q=0$ and $q=3.5$. Moreover, only for the double bound expansion the bands of the form factor at $q=0$ and $q=3.5$ are consistent with each other. The most precise band is the one at $q=0$, which therefore represents our present best estimate of the kaon form factor in the case of the LQCD inputs.
VI.3 Charged kaon radius
@@
where the second error comes from the small dependence upon the choice of the kinematical function (estimated using half of the spread of our results obtained at $q=0$ and $q=3.5$). Thus, the value quoted by PDG underestimates the uncertainty by a factor of $\simeq 2$ due to the model-dependent assumption about the momentum dependence of the charged kaon form factor.
@@ VII Conclusions
We have discussed a specific application of the framework of multiple dispersive bounds, firstly introduced in the companion paper [1], to the study of sub-threshold branch-cuts. The main results are as follows:
  • In the standard BGL $z$-expansion we have proposed a double dispersive bound as the proper way to deal with the presence of sub-threshold branch-cuts in accordance with unitarity. The key idea is that two different unitarity constraints have to be considered at the same time, namely the first one in the the pair-production arc and the second one in the extra region between the lowest sub-threshold and the pair-production branch-points. We have highlighted that the bound related to the extra region, $\chi_{\rm extra}^{U}$, is not related to the usual susceptibilities, i.e. to the quantities coming from the computation of the derivatives of appropriate two-point Green functions in momentum space. It remains an open issue to understand the proper Euclidean correlation functions that would give direct access to the bound $\chi_{\rm extra}^{U}$ directly from first-principles, e.g. from lattice QCD simulations.
@@
  • We have analyzed either the experimental data or the lattice QCD results available for the electromagnetic form factor of the charged kaon in the spacelike-region. The comparison between the results corresponding to the BGL $z$-expansions based either on the single, total or on the double dispersive bound has clearly shown that the latter $z$-expansion provides the most precise extrapolation at large momentum transfer as well as the most stable results with respect to the choice of the outer function outside the pair-production region. We have considered also the predictions of the $z$-expansion based on the Szegő polynomials, proposed in Refs. [26, 27, 28] and adopted also in Refs. [29, 30, 31], showing that the corresponding results for the kaon form factor do not fulfill unitarity on the full circle.
  • We have highlighted the impact of the application of the unitarity filter on the given set of input data (see Section VI.2 and Ref. [1]) and we have compared our model-independent findings for the charged kaon radius with the model-dependent estimates made in the last PDG review [16] and in Ref. [15] (see Section VI.3).